could you help me about this?
Derermine the value (or set of values) of K for which the equation x^2+y^2-2x-4y=K represents a real circle, a point, and an imaginary circle.
i rewrite the equation to (x-1)^2+(y-2)^2=K-5 but i dont know how should i go on =(

Answer by stanbon(75887)

Derermine the value (or set of values) of K for which the equation x^2+y^2-2x-4y=K represents a real circle, a point, and an imaginary circle.
i rewrite the equation to (x-1)^2+(y-2)^2=K-5 but i dont know how should i go on =(
==================
If your work is correct up to that point,
K-5 is the square of the radius of the circle.
-----------------
If sqrt(K-5) is positive you have a real circle.
If sqrt(K-5) is zero you have a point
If sqrt(K-5) is negative you have an imaginary circle.
--------------------------------
Cheers,
Stan H.